# Occlusion-Aware Ground Target Tracking by a Dubins Vehicle Using Visibility Volumes

### v1

@@ 6.4 Controller Turning
is upper bounded for all possible $r\geq\tau R$ and $q_{\psi}\in\mathbb{S}$ because the term inside the parenthesis in (110) is at most $v$, and the term outside the parenthesis is bounded $-\lambda_{\rm max}\leq\frac{\arctan\left(\beta(r-R)\right)}{1+\beta^{2}(r-R)^{2}}\leq\lambda_{\rm max}$ similar to (101). The value of $\lambda_{\rm max}$ was found to occur when $\beta\tilde{r}\arctan(\beta\tilde{r})-\frac{1}{2}=0$ with a value of approximately $\lambda_{\rm max}\approx 0.41195$. The parameter $\beta$ controls the magnitude of both $\dot{\psi}_{d}$ and $u_{\rm lyap}$ which can be tuned such that the controller stays within the limits (2). A grid search for tuning the value of $\beta$ can be seen in Fig. 10 where the POI’s speed is $v_{g}=5$ m/s, the UAV’s speed is $v=20$ m/s, the UAV’s turn radius is $r_{\rm min}=50$ m, the radius of the orbit is $R=85.9$ m, the radius rate of change is $\dot{R}=1.30$ m/s, and the turning parameter is $\beta=0.025$. The maximum value found by the grid search for $r\geq R$ was $|\dot{\psi}_{d}+u_{\rm lyap}|=0.378$ rad/s which is less than the turn rate limit of $\frac{v}{r_{\rm min}}=0.4$ rad/s.
Figure 10: Tuning the $\beta$ parameter pertaining to attracting to the orbit using a grid search. The maximum magnitude of $|\dot{\psi}_{d}+u_{\rm lyap}|$ is shown for all possible headings $q_{\psi}$. Only the points $r\geq 0.8R$ are shown. The shading of the plot indicates the magnitude, ranging from dark to light. The black dashed line shows the radius $R=89.5$ m.
@@ 7 Numerical Study Results
This section presents numerical results that illustrate the reference orbit design strategy of Sec. 5 coupled with the feedback controller of Sec. 6333The implementation of this study can be found at https://github.com/robotics-uncc/DubinsVisibilityTracking. The environment is modeled after the high-rise city-center of Charlotte, North Carolina with building geometry from Open Street Map [66] extracted from a bounding box with corners located at $(80.8471^{\circ}$, $35.2209^{\circ}$) and $(80.8335^{\circ}$, $35.2289^{\circ}$). The POI moves along the road northeast at constant velocity $v_{g}=5$ m/s from $\bm{g}=[-150,-45]^{\rm T}$, then takes a turn and continues southeast along the next road. The POI takes $48.16$ s to reach the end of the second road and the UAV’s motion is simulated over the same period.
Figure 11: The trajectory of the UAV as it tracks a moving POI through an urban environment. The trajectory of the UAV is shown in red as it spirals around the moving POI, attracted to the black visibility orbit constrained within the visibility volumes. The visibility volumes are sliced at a constant altitude and shown in green. The light green dashed outline is the previous visibility volume and the dark green dashed outline is the next visibility volume. The trajectory of the POI is shown in gray.
Two examples will be presented—the first considers a higher altitude UAV with larger turn radius, the second considers a lower altitude UAV with a tighter turn radius.
In the first example, the UAV flies with a constant speed of $v=20$ m/s at an altitude of $h_{\rm UAVUAV}=300$m with a initial position of $\bm{q}=[0,100,-\pi]^{\rm T}$. To track the moving POI, the visibility along the known path is calculated with the adaptive sampling algorithm Alg. 1 with an initial sample spacing of $||\bm{g}_{i}-\bm{g}_{i+}||=20$ m and a cutoff of $d_{\rm cutoff}=20\times 10^{6}$ ${\rm m}^{3}$. The visibility volumes are calculated with $d_{\rm max}=400$ m. The minimum turn radius of the UAV was set to $r_{\rm min}=50$ m, $\beta=0.029$ was selected from Fig. 10, and the steering gain set was to $k_{\psi}=20$. The initial radius of the orbit is $R_{i}=158$ m and the final radius is $R_{f}=127$ m. The trajectory was solved with the Runge-Kutta 45 method [67]. As seen in Fig. 11 the UAV can start at an initial condition off of the visibility orbit with a heading not aligned with the vector fields and steer onto the morphing visibility orbit. The visibility circles from Sec. 5.3 are within the visibility volumes in Fig. 11 and the UAV’s trajectory deviates minimally from the orbits. With the assumption that the POI remains visible during the time steps between the sampled orbits, the UAV can maintain the visibility of the POI for its trajectory.
@@
Figure 12: A plot of the error in the first $20$ s of the trajectory in Fig. 11. The plots on the left are the radial distances $r$ and $R$ and error, while the plots on the right are the headings $\psi_{d}$ and $q_{\psi}$ and the heading error.
@@
The graphs show that the proportional controller with a feed-forward term can force the UAV onto the vector field, which guides the UAV onto the time-varying orbit. The time-varying orbit $O$ is a linear piecewise continuous function with a discontinuous derivative that occurs at the orbit transition points where $\dot{R}$ or $\dot{\bm{g}}$ changes abruptly. This violates the assumption of [54], which assumes a bounded second derivative. The violation of the assumption causes the UAV to deviate from the orbit slightly at each orbit transition point. Still, the controller can quickly force the UAV back onto the orbit. This occurrence can be seen on the error graphs where the error spikes up every time the rate of change of the radius changes. The implementation of the controller considers the minimum of traveling clockwise or counter-clockwise as the heading error, allowing the UAV to turn counter-clockwise from $\pi$ to $-\pi$. The control for the UAV is shown in Fig. 13.
Figure 13: A plot of the heading control effort $u_{\psi}$ (black) for the first $20$ s of the trajectory in Fig. 11. The plot spikes in value at the discontinuities of the orbits’ radial rate of change. The controller is saturated to $u_{\psi,\rm max}=v/r_{\rm min}=0.8$ rad/s (gray dashed) and never exceeds the limits.
There are large spikes in the control effort whenever the rate of change of the radius discontinuously changes. The saturation of the controller only comes into effect in this example when the UAV’s proportional controller exceeds the turn rate control limit 2.
@@
The second example of the tracking approach is shown in Fig. 14. In this example, the UAV has a very sharp turning radius of $r_{\rm min}=5$ m (e.g., representative of a smaller more agile system), a speed of $v=20$ m/s, an altitude of $h_{\rm UAV}=50$ m, an initial configuration of $\bm{q}=[-45,450,0]^{\rm T}$, and a maximum sensing distance of $d_{\rm max}=100$ m. The controller uses $\beta=0.5$ and $k_{\psi}=40$. The visibility volumes were calculated along a different known POI trajectory that starts at $\bm{g}=[-45,430]^{\rm T}$ and moves south west with speed $v_{g}=5$ m/s. As before, Alg. 1 is used to compute the visibility volumes. An initial spacing of $||\bm{g}_{i}-\bm{g}_{i+1}||=20$ m and a cutoff of $d_{\rm cutoff}=500\times 10^{3}$ ${\rm m}^{3}$ are assumed. The more maneuverable UAV makes tighter turns at a lower altitude with stricter sensing constraints allowing for and orbit with an initial radius of $R_{i}=21.5$ m and final radius of $R_{f}=19.7$ m. The UAV orbits the POI for $56.52$ s and circles the POI more times in Fig. 14 compared to Fig. 11. The UAV model quickly turns just over $180^{\circ}$ during the first second of the simulation onto the vector field (76) using (89).
Figure 14: The trajectory of a UAV with a small turn radius as it tracks a moving POI through an urban environment. The trajectory of the UAV is shown in red as it spirals around the moving POI attracted to the black visibility orbit constrained within the visibility volumes. The visibility volumes are sliced at a constant altitude and shown in green. The light green dashed outline is the previous visibility volume and the dark green dashed outline is the next visibility volume. The trajectory of the POI is shown in gray.
During the first $1$s of the trajectory, the control shown in Fig. 15 is saturated to the maximum value given by the turn radius and speed $u_{\psi,\rm max}=v/r_{\rm min}=4$ rad/s because the UAV needs to turn around quickly. The spikes in the controller appear as a result of the discontinuous changes in $\dot{R}$ occurring when the UAV transitions from following the locus of orbits between $R_{i-1}$ and $R_{i}$ to the locus of orbits between $R_{i}$ and $R_{i+1}$. The spikes in Fig. 15 appear smaller than the spikes in Fig. 13 because the maximum control $u_{\psi,\rm max}$ is larger.
Figure 15: A plot of the heading control effort $u_{\psi}$ (black) for the first $20$ s of the trajectory in Fig. 14. The plot spikes in value every at the discontinuities of the orbits’ radial rate of the change. The controller is saturated to $u_{\psi,\rm max}=v/r_{\rm min}=4$ rad/s (gray dashed) and never exceeds the limits.
@@
An example of the error in Fig. 14 is shown in Fig. 16. The plots show the distance of the UAV $r$ to the orbit $O$ in the radial direction and the difference between the desired $\psi_{d}$ and actual heading $\psi$. The initial condition for the UAV starts inside the orbit and grows larger as the UAV turns to minimize the difference between the desired and actual headings. The increased number of orbits about the POI can also be seen in the plot of heading angle $\psi$ by the increase in the number of transitions between $\pi$ and $-\pi$. The increasing nature of the heading angle is because the controller forces counter-clockwise circulation about the POI. To circulate clockwise about the POI used the negative solution of (74).
Figure 16: A plot of the error in the first $20$s of the trajectory in Fig. 14. The plots on the left are the radial distances $r$ and $R$ and error, while the plots on the right are the headings $\psi_{d}$ and $q_{\psi}$ and the heading error.
8 Flight Test
@@
The proposed method was tested on April 27th 2025 using a multicopter UAV to track a moving POI on the University of North Carolina at Charlotte campus. The urban-like environment was modeled using data from OpenStreetMap [66] extracted around the point (35°18’46.3"N 80°44’24.8"W) (see Fig. 17). In the model of the campus, nearby buildings (height 10 m) and trees (height 30 m) may obscure the UAV’s view of the POI. The experiment was performed in an augmented reality setting—a real UAV was flight tested but a virtual gimballed camera was used to track a virtual POI in the environment. The UAV used in the experiment was a custom-built multirotor [68] with an autopilot running Ardupilot [69] and a companion computer for guidance. The path of the POI was obtained prior to the experiment by collecting the GNSS trace of a Clearpath Robotics Jackal unmanned ground vehicle (UGV) steered by a human operator. The UGV’s speed varied from $0.3$ m/s to $0.4$ m/s and was approximated as a couple of constant velocity sections with an instantaneous velocity change at the junction of the two path segments. To create the constant sections, two lines of best fit were fit the corresponding GPS trace segments shown in Fig. 17 resulting in a path with total length of $90.5$ m meters. This pre-recorded path, $\hat{\bm{g}}$, consisting of latitude and longitude positions over time, was then processed on a laptop computer using the adaptive sampling algorithm with an initial sampling of $||\bm{g}_{i}-\bm{g}_{i+1}||=20$ m and a cutoff of $d_{\rm cutoff}=10^{5}$ ${\rm m}^{3}$. The VVs along the path were computed using a maximum sensing radius of $d_{\rm max}=50$ m as shown in Fig. 17. The UAV was assumed to fly at a constant velocity of $v=3$ m/s at an altitude of $h_{\rm UAV}=35$ m and with a minimum turn radius of $r_{\rm min}=5$ m. The altitude was chosen with allowance so that the UAV flies over all the obstacles. The locus of orbits was created with Alg. 2 and can be seen in Fig. 17. The orbits start large because the UAV is far from the trees and buildings (approximately 7.8 m from the building and 24.3 m from the trees), and then shrink once the UGV turns the corner because the UGV is closer to the building (approximately 3.2 m from the building and 34.3 m from the trees).
Figure 17: Left: Three-dimensional model of the campus of the University of North Carolina at Charlotte, where the experiment will occur. The buildings are shaded red, and the trees are shaded green. The cyan path is the approximated POI’s path. The VVs are shown in transparent white. Right: A satellite view of the POI path on campus. The POI’s path is red, and the estimated path is black. The visibility orbits are shown as the purple to yellow circles.
To control the UAV, the onboard companion computer uses the Robot Operating System (ROS) Noetic [70] to communicate with the UAV. Since the experiment was conducted with a multirotor, rather than a fixed-wing vehicle, the heading of the vehicle could be independently controlled from the velocity. As a result, the velocity from the designed vector field (78) (with gain $\beta=0.5$) was used directly as a setpoint command to the autopilot via a ROS node. The required POI trajectory data to execute the guidance and control strategy is the set of orbit radii, orbit origin positions, and corresponding times.
@@
To minimize the degree to which the virtual gimballed camera must actuate (since real gimbals have actuation limits), the UAV’s yaw is controlled so the relative angle to the virtual POI is consistent with the guidance vector field (which tends to face the future position of the POI). The same ROS node that controls the velocity controls the yaw rate using a proportional controller with feed-forward term,
$\dot{\psi}=k_{\rm yaw}(\psi_{d}-\psi)+\dot{\psi}_{d}\;,$
where the UAV’s yaw is denoted $\psi$, $k_{\rm yaw}=0.6$ is a positive yaw gain, the desired yaw is (79), and the desired yaw rate is (88).
@@
A software-in-the-loop (SITL) simulation was created to test experimental parameters before flying an experiment. The simulation used ROS with Gazebo Classic and the Ardupilot SITL package. The simulation was used to test different sensing constraints and altitudes to find a set of parameters for the flight experiment. The SITL simulation was used to compare the proposed approach of using a time-varying radius $R$ from Sec. 5 to a naive approach of orbiting with a constant radius. The time-varying radius $R$ had a minimum value of $11.2$ m and a maximum value of $23.4$ m. Moreover, given the assumed maximum sensing distance of $d_{\rm max}=50$ m and the UAV’s altitude $h_{\rm UAV}=35$ m the maximum allowable orbit radius (in the absence of obstacles) is $35$ m. Therefore, the three constant-radii SITL simulations used for comparison: $11.2$ m, $23.4$ m, and $35.0$ m.
@@
The visibility of the virtual POI during each SITL simulation is shown in Fig. 18. The line between the simulated POI and UAV was ray-traced after the mission completed to test for visibility. From Fig. 18, the visibility of the POI decreases as the fixed radius increases. However, the dynamic radius maintains $99\%$ visibility while maximizing the standoff distance.
Figure 18: A bar plot showing the percentage of time the POI is visible from the UAV during a simulated mission.
An image of the experiment setup is shown in Fig. 19.
Figure 19: Aerial images showing the UAV and its environment. The UAV maneuvers around the building and trees while the virtual target follows the orange line.
@@
In the real-world flight experiment, the virtual POI moves along the cyan path in Fig. 17 and the orange path in Fig. 19, while the UAV circles above. The UAV hovers above the virtual POI until the mission is started, and both the virtual POI and real UAV start moving at the same time. A total of six flight tests were conducted: three using the time-varying, visibility-informed $R$ orbits and three using a constant-radius orbit of $R=35$ m to highlight the importance of considering occlusions. The resulting visibility percentage of the POI during each experiment are shown in Fig. 20 along with signed radial error statistics (computed after the UAV converged to the orbit). The visibility was found in post-processing by ray-tracing the line between the real UAV position and the virtual POI position in the virtual environment. The UAV was able to successfully view the POI an average of 98.3% of the time using the proposed method in the experiment; This was on average 29% more than the fixed non-visibility method for the same tracking mission. The overall visibility percentages are comparable to simulation results in Fig. 18. The proposed method did not produce 100% visibility due to tracking errors.
Figure 20: Right: A bar plot showing the percentage of time during the mission when the POI is visible from the UAV. Left: A violin plot showing the orbit tracking error during the experiments. The width of the violin plots shows the density of the errors. The horizontal bars show the minimum and maximum errors during the experiments. The black circle in the middle shows the mean value.
The radial error statistics in the right panel of Figure 20 have a non-zero mean in both cases. This can be explained by the controller sending fixed velocity setpoints at 10 Hz based on the estimated position rather than the continuous control signal in Sec. 7. This results in a tendency to err on the outside, rather than inside, of the orbit. The higher error values in the visibility-informed orbit method are not surprising since a time-varying radius with tighter turn radii is more difficult to track.
Figure 21 provides a more detailed time-history of the trajectory and radial error for one of the experiments using the visibility-informed orbit.
Figure 21: The UAV and the POI trajectories during a visibility-informed orbit experiment. Left: The trajectories overlayed on satellite images. The red trajectory is the POI’s and the black trajectory is the UAV’s. Selected images of the virtual POI are shown. The orange lines represent a visible POI (dark blue cube), while the blue lines represent an obscured POI. Right: The radial error of the UAV to the planned circular orbit and the visibility of the POI from the UAV’s camera.
@@
The left panel of Figure 21 shows the trajectory of the UAV continuously circling the virtual POI. Selected images from the virtual UAV’s camera sensor are shown. Two of the images show the center point of the POI visible (rendered as a dark blue cube resting on top of a cyan car), while another shows the center point of the POI obscured by the corner of the red building. The radial error in the right panel of Fig. 21 shows minimal deviation from the circular orbit during the experiment.
@@
Lastly, Fig. 22 shows results from one the experiments using the fixed orbit radius wherein the visibility was 66.5%.
Figure 22: The trajectory of the UAV and the POI during the non-visibility-informed mission. Left: The trajectories overlayed on satellite images. The red trajectory is the POI’s and the black trajectory is the UAV’s. Selected images of the virtual POI are shown. The orange lines represent a visible POI (dark blue cube), while the blue lines represent an obscured POI. Right: The radial error of the UAV to the planned circular orbit and the visibility of the POI from the UAV’s camera.
Two of the images in the left panel of Fig. 22 show the POI obscured by the trees (green) and building (red), respectively. The radial error on the constant radius orbit is lower than along the time-varying radius (reflecting the same results observed in simulation).

### v2

@@ 5.4 Controller Turning
is upper bounded for all possible $r\geq\tau R$ and $q_{\psi}\in\mathbb{S}$ because the term inside the parenthesis in (106) is at most $v$, and the term outside the parenthesis is bounded $-\lambda_{\rm max}\leq\frac{\arctan\left(\beta(r-R)\right)}{1+\beta^{2}(r-R)^{2}}\leq\lambda_{\rm max}$ similar to (98). The approximate maximum is $\lambda_{\rm max}\approx 0.41195$ and occurs when $\beta\tilde{r}\arctan(\beta\tilde{r})-\frac{1}{2}=0$. The parameter $\beta$ controls the magnitude of both $\dot{\psi}_{d}$ and $u_{\rm lyap}$, and is tuned so the controller stays within the limits (2). A grid search the value of $\beta$ can be seen in Fig. 9 where the POI’s speed is $v_{g}=5$ m/s, the UAV’s speed is $v=20$ m/s, the UAV’s turn radius is $r_{\rm min}=50$ m, the radius of the orbit is $R=85.9$ m, the radius rate of change is $\dot{R}=1.30$ m/s, and the tuning parameter $\beta=0.025$. The maximum value found by the grid search for $r\geq R$ was $|\dot{\psi}_{d}+u_{\rm lyap}|=0.378$ rad/s which is less than the turn rate limit of $\frac{v}{r_{\rm min}}=0.4$ rad/s.
Figure 9: Tuning the $\beta$ parameter pertaining to attracting to the orbit using a grid search.
@@ 6 Numerical Study Results
This section presents numerical results from the reference orbit design strategy of Sec. 4 and feedback controller of Sec. 5333The implementation of this study can be found at https://github.com/robotics-uncc/DubinsVisibilityTracking. The environment is modeled after the city-center of Charlotte, North Carolina with building geometry from Open Street Map [29] extracted from a bounding box, $(80.8471^{\circ}$, $35.2209^{\circ}$) and $(80.8335^{\circ}$, $35.2289^{\circ}$). The POI moves along the road northeast at constant velocity $v_{g}=5$ m/s from $\bm{g}=[-150,-45]^{\rm T}$, then takes a turn and continues southeast along the next road. The POI takes $48.16$ s to reach the end of the second road and the UAV’s motion is simulated over the same period.
Figure 10: Two examples of a UAV (red) tracking a moving POI through an urban environment (gray).
Two examples are presented in Fig. 10—the first and top considers a higher altitude UAV with a larger turn radius, the second and bottom considers a lower altitude UAV with a tighter turn radius.
In the first example, the UAV flies with a constant speed of $v=20$ m/s at an altitude of $h_{\rm UAV}=300$m with a initial position of $\bm{q}=[0,100,-\pi]^{\rm T}$. To track the moving POI, the visibility along the known path is calculated with the adaptive discretization algorithm, Sec. 4.2, with an initial sample spacing of $||\bm{g}_{i}-\bm{g}_{i+}||=20$ m and a cutoff of $d_{\rm cutoff}=20\times 10^{6}$ ${\rm m}^{3}$. The VVs are calculated with $d_{\rm max}=400$ m. The minimum turn radius of the UAV was set to $r_{\rm min}=50$ m, $\beta=0.025$ was selected from Fig. 9, and the steering gain set was to $k_{\psi}=20$. The initial radius of the orbit is $R_{i}=158$ m and the final radius is $R_{f}=127$ m. The trajectory was solved with the Runge-Kutta 45 method [30]. As seen in the two trajectories in Fig. 10 the UAV (red trajectories) can start at initial conditions off of the visibility orbit with a heading not aligned with the vector fields and steer onto the morphing visibility orbits. The visibility orbits (black dashed circles) from Sec. 4.3 are within the VVs (dashed green lines) in Fig. 10, and the UAV’s trajectories deviate minimally from the orbits. With the assumption that the POI remains visible during the time steps between the sampled orbits, the UAV can maintain the visibility of the POI for its trajectory (dashed gray).
@@
Figure 11: A plot of the error and control in the first $20$ s of the simulations (first: left and second: right).
@@
The graphs show that the proportional controller with a feed-forward term forces the UAV onto the vector field, which guides the UAV onto the time-varying orbit. The time-varying orbit $O$ is a linear piecewise continuous function with a discontinuous derivative that occurs at the orbit transition points where $\dot{\bm{g}}$ changes abruptly. This violates the assumption of [26], which assumes a bounded second derivative. The violation of the assumption causes the UAV to deviate from the orbit slightly at each orbit transition point. Still, the controller can quickly force the UAV back onto the orbit. This occurrence can be seen on the error graphs where the error spikes up every time the rate of change of the radius changes. The implementation of the controller considers the minimum of traveling clockwise or counter-clockwise as the heading error, allowing the UAV to turn counter-clockwise from $\pi$ to $-\pi$. The control for the UAV is shown in Fig. 11. There are large spikes in the control effort whenever the rate of change of the radius discontinuously changes. The saturation of the controller only comes into effect in this example when the UAV’s proportional controller exceeds the turn rate control limit (2).
@@
The second example of the tracking approach is shown on the bottom of Fig. 10. In this example, the UAV has a very sharp turning radius of $r_{\rm min}=5$ m (e.g., representative of a smaller more agile system), a speed of $v=20$ m/s, an altitude of $h_{\rm UAV}=50$ m, an initial configuration of $\bm{q}=[-45,450,0]^{\rm T}$, and a maximum sensing distance of $d_{\rm max}=100$ m. The controller uses $\beta=0.5$ and $k_{\psi}=40$. The VVs were calculated along a different known POI trajectory that starts at $\bm{g}=[-45,430]^{\rm T}$ and moves southwest with speed $v_{g}=5$ m/s. As before, the adaptive discretization algorithm in Sec. 4.2 is used to compute the VVs. An initial spacing of $||\bm{g}_{i}-\bm{g}_{i+1}||=20$ m and a cutoff of $d_{\rm cutoff}=500\times 10^{3}$ ${\rm m}^{3}$ are assumed. The more maneuverable UAV makes tighter turns at a lower altitude with stricter sensing constraints allowing for and orbit with an initial radius of $R_{i}=21.5$ m and final radius of $R_{f}=19.7$ m. The UAV orbits the POI for $56.52$ s and circles the POI more times in bottom of Fig. 10 compared to the top Fig. 10. The trajectory of the UAV is shown in red, and the attractive VO is shown in black. The UAV model quickly turns just over $180^{\circ}$ during the first second of the simulation onto the vector field (74) using (86). The UAV is able to remain on the VO inside the VPs (green dashed), maintaining a view of the POI (gray dashed). During the first $1$s of the trajectory, the control shown in Fig. 11 is saturated to the maximum value given by the turn radius and speed $u_{\psi,\rm max}=v/r_{\rm min}=4$ rad/s because the UAV needs to turn around quickly. The spikes in the controller appear as a result of the discontinuous changes in $\dot{R}$ occurring when the UAV transitions from following the locus of orbits between $R_{i-1}$ and $R_{i}$ to the locus of orbits between $R_{i}$ and $R_{i+1}$. The spikes in the plot of the control for the second simulation in Fig. 11 appear smaller than the spikes in the plot of the control for the first simulation in Fig. 11 because the maximum control $u_{\psi,\rm max}$ is larger.
@@
An example of the error for the second simulation is shown on the right half of Fig. 11. The plots show the distance of the UAV $r$ to the orbit $O$ in the radial direction and the difference between the desired $\psi_{d}$ and actual heading $\psi$. The initial condition for the UAV starts inside the orbit and grows larger as the UAV turns to minimize the difference between the desired and actual headings. The increased number of orbits about the POI can also be seen in the plot of heading angle $\psi$ by the increase in the number of transitions between $\pi$ and $-\pi$. The increasing nature of the heading angle is because the controller forces CCW circulation about the POI. To circulate CW, use the negative solution of (72).
7 Flight Test Experiment
@@
The proposed method was tested on April 27th, 2025, using a multicopter UAV to track a moving POI on the University of North Carolina at Charlotte campus. The urban-like environment was modeled using data from OpenStreetMap [29] extracted around the point (35°18’46.3"N 80°44’24.8"W) (see Fig. 12). In the model of the campus, nearby buildings (height 10 m) and trees (height 30 m) may obscure the UAV’s view of the POI. The experiment was performed in an augmented reality setting—a real UAV was flight tested, but a virtual gimballed camera was used to track a virtual POI in the environment. The UAV used in the experiment was a custom-built multirotor [31] with an autopilot running Ardupilot [32] and a companion computer for guidance. The path of the POI was obtained prior to the experiment by collecting the GNSS trace of a Clearpath Robotics Jackal unmanned ground vehicle (UGV) steered by a human operator. The UGV’s speed varied from $0.3$ m/s to $0.4$ m/s and was approximated as a couple of constant velocity sections with an instantaneous velocity change at the junction of the two path segments. To create the constant sections, two lines of best fit (cyan lines) were fitted to the corresponding GPS trace segments Fig. 12, resulting in a path with a total length of $90.5$ m. This pre-recorded path, $\hat{\bm{g}}$, consisting of latitude and longitude positions over time, was then processed on a laptop computer using the adaptive discretization algorithm with an initial discretization of $||\bm{g}_{i}-\bm{g}_{i+1}||=20$ m and a cutoff of $d_{\rm cutoff}=10^{5}$ ${\rm m}^{3}$. The VVs along the path were computed using a maximum sensing radius of $d_{\rm max}=50$ m shown as white volumes in Fig. 12. The UAV was assumed to fly at a constant velocity of $v=3$ m/s at an altitude of $h_{\rm UAV}=35$ m and with a minimum turn radius of $r_{\rm min}=5$ m. The altitude was chosen with allowance so that the UAV flies over all the obstacles. The locus of orbits was created with Alg. 1. To control the UAV, the onboard companion computer uses the Robot Operating System (ROS) Noetic [33] to communicate with the UAV. Since the experiment was conducted with a multirotor, rather than a fixed-wing vehicle, the heading of the vehicle could be independently controlled from the velocity. As a result, the velocity from the designed vector field (76) (with gain $\beta=0.5$) was used directly as a setpoint command to the autopilot via a ROS node. The required POI trajectory data to execute the guidance and control strategy is the set of orbit radii, orbit origin positions, and corresponding times.
@@
To minimize the degree to which the virtual gimballed camera must actuate (since real gimbals have actuation limits), the UAV’s yaw is controlled so the relative angle to the virtual POI is consistent with the guidance vector field (which tends to face the future position of the POI). The same ROS node that controls the velocity controls the yaw rate using a proportional controller with feed-forward term, $\dot{\psi}=k_{\rm yaw}(\psi_{d}-\psi)+\dot{\psi}_{d}\;,$ where the UAV’s yaw is denoted $\psi$, $k_{\rm yaw}=0.6$ is a positive yaw gain, the desired yaw is (77), and the desired yaw rate is (85).
@@
A software-in-the-loop (SITL) simulation was created to test before flying an experiment. The simulation used ROS with Gazebo Classic and the Ardupilot SITL package. To test different sensing constraints and altitudes to find a set of parameters for the flight experiment. The SITL simulation compared the proposed approach of using a time-varying radius $R$ from Sec. 4 to orbiting with a constant radius. The time-varying radius $R$ had a minimum value of $11.2$ m and a maximum value of $23.4$ m. Moreover, given the assumed maximum sensing distance of $d_{\rm max}=50$ m and the UAV’s altitude $h_{\rm UAV}=35$ m the maximum allowable orbit radius (in the absence of obstacles) is $35$ m. Therefore, the three constant-radii SITL simulations used for comparison: $11.2$ m, $23.4$ m, and $35.0$ m.
@@
The visibility of the virtual POI during each SITL simulation was 100%, 84.0%, 64.0%, and 99.3% for orbit radii of 11 m, 23 m, 35 m, and the dynamic radius $R(t)$, respectively. The POI’s visibility from UAV was tested using ray-tracing after the mission. From the simulations, the POI’s visibility decreases as the radius increases. However, the dynamic radius maintains $99.3\%$ visibility while maximizing the standoff distance. An image of the experiment setup is shown in Fig. 12.
Figure 12: Model of the simulated and real-world environment used during the experiment.
@@
In the real-world flight experiment, the virtual POI moves along the cyan path in Fig. 12, while the UAV circles above. The UAV hovers above the virtual POI until the mission is started, and both the virtual POI and real UAV start moving at the same time. A total of six flight tests were conducted: three using the time-varying, visibility-informed $R$ orbits and three using a constant-radius orbit of $R=35$ m to highlight the importance of considering occlusions. The resulting visibility percentage of the POI during each experiment is 98.0%, 98.0%, and 99.0% for the VOs and 70.3%, 71.1%, and 66.5% for the fixed non-visibility orbits. The visibility was found in post-processing by ray-tracing the line between the real UAV position and the virtual POI position in the virtual environment. The UAV was able to successfully view the POI an average of 98.3% of the time using the proposed method in the experiment; This was on average 29% more than the fixed non-visibility method for the same tracking mission. The overall visibility percentages are comparable to the SITL simulations. The proposed method did not produce 100% visibility due to tracking errors. Figure 13 provides a more detailed time-history of the trajectory and radial error for one of the experiments using the visibility-informed orbit.
Figure 13: The trajectory, visibility, and radial error of the UAV and the POI for a visibility-informed orbit.
@@
The left panel of Fig. 13 shows the trajectory of the UAV continuously circling the virtual POI. Selected images from the virtual UAV’s camera sensor are shown. Two of the images show the center point of the POI visible (rendered as a dark blue cube resting on top of a cyan car), while another shows the center point of the POI obscured by the corner of the red building. The VO, show as gray dashed lines, start large because the UAV is far from the trees and buildings (approximately 7.8 m from the building and 24.3 m from the trees), and then shrink once the UGV turns the corner because the UGV is closer to the building (approximately 3.2 m from the building and 34.3 m from the trees). The radial error in the right panel of Fig. 13 shows a mean deviation of $0.72$ m with a maximum of $1.77$ m and a minimum of $-0.23$ m after converging to the circular orbit during the experiment.
@@
Lastly, Fig. 14 shows results from an experiment using the fixed orbit radius wherein the visibility was 66.5%.
Figure 14: The trajectory, visibility, and radial error of the UAV and the POI for a non-visibility-informed orbit.
Two of the images in the left panel of Fig. 14 show the POI obscured by the trees (green) and building (red). The mean radial error after converging to the constant radius orbit is $0.29$ m with a maximum value of $0.81$ m and a minimum value of $-0.22$ m. Both experiments have non-zero mean radial error. This can be explained by the controller sending fixed velocity setpoints at 10 Hz based on the estimated position rather than the continuous control signal in Sec. 6. This results in a tendency to err on the outside, rather than inside, of the orbit.